OpenStudy (anonymous):

related rates problem: A street light is mounted at the top of a 15ft pole. A 6ft tall man walks away from the pole at 5ft/second along a straight path. How fast is the top of his shadow moving when he is 40ft away from the pole?

OpenStudy (amistre64):

might need a right triangle for this

OpenStudy (anonymous):

yup i have an answer but Im not sure if i achieved it the correct way

OpenStudy (amistre64):



OpenStudy (amistre64):

we have similar triangles: such that: \[\frac{15}{w+s}=\frac{6}{s}\]

OpenStudy (amistre64):

we know the rate of w with repect to time; w' = 5

OpenStudy (amistre64):

s' = w' + s' in this case I beleive

OpenStudy (amistre64):

15s = 6w+6s 15s-6s = 6w s(15-6) = 6w s = 6w/9 = 2w/3 ds/dw = 2w'/3

OpenStudy (amistre64):

w' = 5, s' = 10/3 if i did it right so w'+s' = 25/3 any idea if thats good?

OpenStudy (anonymous):

thats what i got haha. I'm studying your explanation more thoroughly...

OpenStudy (amistre64):

im always ;eary of this one casue the 40 feet seems to be superfluous to the overall mathing

OpenStudy (amistre64):

leary even lol

OpenStudy (amistre64):

example 6 http://tutorial.math.lamar.edu/Classes/CalcI/RelatedRates.aspx

OpenStudy (anonymous):

many thanks that is just what i needed